# Simulations of inertial liquid-lens coalescence with the pseudopotential lattice Boltzmann method

### v1

@@ I Introduction
The coalescence of droplets is a fundamental process relevant to natural phenomena and many industrial applications, such as formation of rain drops Low1982, stability of emulsions Goff1997; Kumar1996, enhanced oil recovery Perazzo2018, coating Eslamian2018 and printing Wijshoff2018. Coalescence is initiated when two droplets come into contact and form a growing liquid bridge. Finally, the two droplets merge into a single droplet, which relaxes to its equilibrium shape. The growth of the bridge is controlled by the interplay of capillarity, viscosity and inertia, and exhibits three different dynamic regimes: the inertial limited viscous regime Paulsen2012; Paulsen2013; Anthony2020, the viscous regime Hopper1984; Eggers1999; Hopper1990, and the inertial regime Aarts2005; Paulsen2011. The coalescence of suspended droplets in a fluid phase Eggers1999; Duchemin2003; Aarts2005; Thoroddsen2007; Paulsen2011; Anthony2020 and sessile droplets on a substrate Ristenpart2006; Narhe2008; Lee2012; HernandezSanchez2012; Eddi2013; Stone2022 have been investigated extensively. At the initial stage of coalescence, in the viscous regime, the radius of the bridge grows as $r_{0}\sim t$ for both spherical droplets and sessile droplets HernandezSanchez2012, whereas, in the inertial regime, the radius of bridge grows following $r_{0}\sim t^{1/2}$ for spherical droplets Burton2007; Paulsen2011; Xia2019 and $r_{0}\sim t^{2/3}$ for sessile droplets on a substrate Eddi2013, which indicates that the geometry of droplets has a strong effect on the coalescence dynamics. The crossover from the viscous regime to the inertial regime of the bridge growth for spherical drops depends on fluid properties and drop size Paulsen2011; Xia2019.
@@ II Method
where $\omega_{i}$ is a coefficient depending on the direction: $\omega_{0}=1/3$ for the zero velocity, $\omega_{1,\dots,6}=1/18$ for the six nearest neighbors and $\omega_{7,\dots,18}=1/36$ for the nearest neighbors in diagonal direction. $c_{s}=\frac{1}{\sqrt{3}}\frac{\Delta x}{\Delta t}$ is the speed of sound. The macroscopic variables, densities and velocities are updated as $\rho^{c}(\vec{x},t)=\rho_{0}\sum_{i}f^{c}_{i}(\vec{x},t)$, where $\rho_{0}$ is a reference density, and $\vec{u}^{c}(\vec{x},t)=\sum_{i}f^{c}_{i}(\vec{x},t)\vec{c}_{i}/\rho^{c}(\vec{x},t)$, respectively. Lattice Boltzmann method can be treated as an alternative solver of Navier-Stokes equation in the limit of small Knnudsen and Mach numbers Succi2001. In our simulations, we choose the lattice constant $\Delta x$, the timestep $\Delta t$, the reference density $\rho_{0}$ and the relaxation time $\tau^{c}$ to be unity, which results in a kinematic viscosity $\nu^{c}$ == $\frac{1}{6}$ in lattice units.
@@
in which $\vec{x}^{\prime}$ denote the nearest neighbours of lattice site $\vec{x}$ and $g_{cc^{\prime}}$ is a coupling constant determining the surface tension. $\Psi^{c}(\mathbf{x},t)$ is called “effective mass”, defined as
@@
The force $\vec{F}_{\mathrm{C}}^{c}(\vec{x},t)$ is then incorporated into the lattice Boltzmann equations by adding a shift $\Delta\vec{u}^{c}(\vec{x},t)=\frac{\tau^{c}\vec{F}_{\mathrm{C}}^{c}(\vec{x},t)}{\rho^{c}(\vec{x},t)}$ to the velocity $\vec{u}^{c}(\vec{x},t)$ in the equilibrium distribution function $f_{i}^{\mathrm{eq}}$. The pseudopotential multicomponent LB method is a diffuse interface method with an interface width of $\approx 5\Delta x$, which elegantly removes the stress singularity at the moving contact line usually occurring in sharp-interface models.
@@ III.2 Two liquid lenses - 2D
We conduct simulations of liquid lens coalescence in two dimensions. To capture the initial stage of coalescence where the bridge growth is not affected by the finite height of the lens, a sufficiently large initial height of the lens is required. We initialize a spherical cap-shaped lens at the fluid interface with maximal height $H=1000$ and contact angles $\theta=22^{\circ},35^{\circ},59^{\circ}$, respectively. After reaching equilibrium, we numerically mirror the lens and locate it at a distance of $\sim 2$ lattice nodes from the original lens. We note that the initial distance between two lenses has no significant influence on coalescence dynamics Hessling2017, and coalescence is predominantly driven by surface tensions.
@@
We show the time evolution of the vertical height of the bridge center $h_{0}$ in Fig. 6(a) for different contact angles. The bridge grows faster with increasing contact angle. For small contact angles $\theta<40^{\circ}$, the simulation results (symbols) agree quantitatively with the theoretical analysis Eq. (15) (solid lines). For a large contact angle $\theta=59^{\circ}$, the similarity solution based on the thin-sheet equations overestimates the bridge growth. We note that our LBM method solves the Navier-Stokes equations and is valid for the whole range of contact angles. In the case of a contact angle of $\theta=90^{\circ}$, we simulate the coalescence of two suspended spherical droplets in another liquid, and interestingly, we find that the theoretical prediction based on the thin-sheet equations underestimates the bridge growth - contrary to that at a contact angle of $\theta=59^{\circ}$.
@@ III.3 Two liquid lenses - 3D
In Fig. 8(a), we show the time evolution of the bridge height $h_{0}$ and radius $r_{0}$ for different contact angles $\theta=60^{\circ}$, $43^{\circ}$, respectively. The vertical height of the bridge grows faster with a larger contact angle, similar to what was observed in the $2D$ simulations (see Fig. 6(a)). Interestingly, the horizontal radius of the bridge grows at the same speed for both contact angles in the initial stage, which indicates that the bridge radius growth is not affected by the initial contact angle of the lenses, and the growth of the bridge radius is not a linear function of the growth of bridge height, in contrast to that observed in coalescing sessile droplets on a substrate. A possible explanation is that due to the no-slip boundary condition, the coalescence speed of sessile droplets on a substrate is relatively slow, giving sufficient time for the cross-section of the bridge to relax immediately to a spherical cap shape with a corresponding equilibrium contact angle. However, the coalescence speed of liquid lenses at a fluid-fluid interface is significantly faster, and the bridge height and radius grow in a non-coupled manner, resulting in a non-linear dependence of bridge radius and height growth.
@@ IV Conclusion
We numerically investigate the inertial coalescence of liquid lenses over a wide range of contact angles using the pseudopotential lattice Boltzmann method. Our simulations in two dimensions successfully capture the self-similar dynamics of both the bridge and velocity profiles, showing quantitative agreement with experimental observations and a theoretical framework based on thin-sheet equations. Furthermore, our simulation results demonstrate that the thin-sheet equations are applicable to describe the quantitative behavior of bridge growth for small contact angles approximately up to $\theta<40^{\circ}$. In the three-dimensional case, we find out that the bridge height grows faster with a larger contact angle, while the growth of the bridge radius is independent of the contact angle at the initial stage. The contact angle of the cross-section of the bridge increases rapidly in the initial stage. It reaches a plateau at the intermediate stage, indicating a transition from a non-linear to a linear dependency between the growth of bridge height and bridge width.
The pseudopotential lattice Boltzmann method has been widely applied to study binary fluid component systems Timm2016; liu_multiphase_2016; xie_effect_2025, and our results demonstrate its applicability to ternary fluid systems, potentially inspiring its use for investigating printing and coating multi-component solutions for functional material synthesis in catalytic and electronic applications vinodh_recent_2024; steinberger_challenges_2024.

### v2

@@ Abstract
Coalescence, liquid lenses, lattice Boltzmann method
@@ I Introduction
The coalescence of droplets is a fundamental process relevant to natural phenomena and many industrial applications, such as formation of rain drops Low1982, stability of emulsions Goff1997; Kumar1996, enhanced oil recovery Perazzo2018, coating Eslamian2018 and printing Wijshoff2018. Coalescence is initiated when two droplets come into contact and form a growing liquid bridge. Finally, the two droplets merge into a single droplet, which relaxes to its equilibrium shape. The growth of the bridge is controlled by the interplay of capillarity, viscosity and inertia, and exhibits three different dynamic regimes: the inertial limited viscous regime Paulsen2012; Paulsen2013; Anthony2020, the viscous regime Hopper1984; Eggers1999; Hopper1990, and the inertial regime Aarts2005; Paulsen2011. These regimes are typically characterized by the Ohnesorge number, $Oh=\eta/\sqrt{\rho R\gamma}$, where $\eta$ is the dynamic viscosity, $\rho$ is the droplet liquid density, $R$ is the droplet radius, and $\gamma$ is the surface tension. This dimensionless number measures the dominance of viscosity over inertia. For $Oh\gg 1$, coalescence occurs in the viscous regime, whereas $Oh\ll 1$ indicates the inertial regime. The corresponding characteristic timescales are the viscous time, $\tau_{v}=\eta R/\gamma$, and the inertial time, $\tau_{i}=\sqrt{\rho R^{3}/\gamma}$ Jacco2025. The coalescence of suspended droplets in a fluid phase Eggers1999; Duchemin2003; Aarts2005; Thoroddsen2007; Paulsen2011; Anthony2020 and sessile droplets on a substrate Ristenpart2006; Narhe2008; Lee2012; HernandezSanchez2012; Eddi2013; Stone2022 have been investigated extensively. At the initial stage of coalescence, in the viscous regime, the radius of the bridge grows as $r_{0}\sim t$ for both spherical droplets and sessile droplets HernandezSanchez2012, whereas, in the inertial regime, the radius of bridge grows following $r_{0}\sim t^{1/2}$ for spherical droplets Burton2007; Paulsen2011; Xia2019 and $r_{0}\sim t^{2/3}$ for sessile droplets on a substrate Eddi2013, which indicates that the geometry of droplets has a strong effect on the coalescence dynamics. The crossover from the viscous regime to the inertial regime of the bridge growth for spherical drops depends on fluid properties and drop size Paulsen2011; Xia2019.
@@ II Method
where $\omega_{i}$ is a coefficient depending on the direction: $\omega_{0}=1/3$ for the zero velocity, $\omega_{1,\dots,6}=1/18$ for the six nearest neighbors and $\omega_{7,\dots,18}=1/36$ for the nearest neighbors in diagonal direction. $c_{s}=\frac{1}{\sqrt{3}}\frac{\Delta x}{\Delta t}$ is the speed of sound. The macroscopic variables, densities and velocities are updated as $\rho^{c}(\vec{x},t)=\rho_{0}\sum_{i}f^{c}_{i}(\vec{x},t)$, where $\rho_{0}$ is a reference density, and $\vec{u}^{c}(\vec{x},t)=\sum_{i}f^{c}_{i}(\vec{x},t)\vec{c}_{i}/\rho^{c}(\vec{x},t)$, respectively. The single relaxation time (SRT) scheme is employed in this work. While the multi-relaxation time (MRT) scheme Fan2010 can enhance stability for high Reynolds number or low-viscosity flows, it incurs a higher computational cost due to the required moment-space transformations. Given the low Reynolds numbers and moderate viscosities considered here, SRT provides a stable and computationally efficient alternative.
Lattice Boltzmann method can be treated as an alternative solver of Navier-Stokes equation in the limit of small Knnudsen and Mach numbers Succi2001. In our simulations, we choose the lattice constant $\Delta x$, the timestep $\Delta t$, the reference density $\rho_{0}$ and the relaxation time $\tau^{c}$ to be unity, which results in a kinematic viscosity $\nu^{c}$ == $\frac{1}{6}$ in lattice units.
@@
in which $\vec{x}^{\prime}$ denote the nearest neighbours of lattice site $\vec{x}$ and $g_{cc^{\prime}}$ is a coupling constant determining the surface tension. $\Psi^{c}(\mathbf{x},t)$ is called “effective mass”, defined as Shan1993; Shan1994
@@
This effective mass formulation ensures that at low densities it approximates the density $\rho^{c}$ itself, while at high densities it asymptotically approaches a saturation limit CHEN2014210. The saturation behavior prevents the collapse of the high-density phase, thereby enhancing numerical stability in simulations.
The force $\vec{F}_{\mathrm{C}}^{c}(\vec{x},t)$ is then incorporated into the lattice Boltzmann equations by adding a shift $\Delta\vec{u}^{c}(\vec{x},t)=\frac{\tau^{c}\vec{F}_{\mathrm{C}}^{c}(\vec{x},t)}{\rho^{c}(\vec{x},t)}$ to the velocity $\vec{u}^{c}(\vec{x},t)$ in the equilibrium distribution function $f_{i}^{\mathrm{eq}}$. The pseudopotential multicomponent LB method is a diffuse interface method with an interface width of $\approx 5\Delta x$, which elegantly removes the stress singularity at the moving contact line usually occurring in sharp-interface models. We note that in the pseudopotential lattice Boltzmann method, the surface tension is governed by the interaction parameter $g_{cc^{\prime}}$ in Eq. (4) and is calculated by the Young-Laplace equation. For a spherical droplet of radius $R$ immersed in another fluid, the Young–Laplace equation, $\gamma=\frac{R\Delta P}{2}$, relates the pressure difference $\Delta P$ over the interface between two fluids to the surface tension $\gamma$.
@@ III.2 Two liquid lenses - 2D
We conduct simulations of liquid lens coalescence in two dimensions. To capture the initial stage of coalescence where the bridge growth is not affected by the finite height of the lens, a sufficiently large initial height of the lens is required. We initialize a spherical cap-shaped lens at the fluid interface with maximal height $H=1000$ and contact angles $\theta=22^{\circ},35^{\circ},59^{\circ}$, respectively. We note that the contact angles were measured after the initially configured single-droplet lens had relaxed to equilibrium. After reaching equilibrium, we numerically mirror the lens and locate it at a distance of $\sim 2$ lattice nodes from the original lens. We note that the initial distance between two lenses has no significant influence on coalescence dynamics Hessling2017, and coalescence is predominantly driven by surface tensions. Given that the lenses approach with negligible relative velocity $v$, the associated Weber number $We=\frac{\rho v^{2}}{\gamma R}$ is much less than $1$, which favors coalescence over phenomena like bouncing or breakup. The surface tension during coalescence remains constant, as the system is both isothermal and free of surfactants.
@@
We show the time evolution of the vertical height of the bridge center $h_{0}$ in Fig. 6(a) for different contact angles. The bridge grows faster with increasing contact angle. For small contact angles $\theta<40^{\circ}$, the simulation results (symbols) agree quantitatively with the theoretical analysis Eq. (15) (solid lines). For a large contact angle $\theta=59^{\circ}$, the similarity solution based on the thin-sheet equations overestimates the bridge growth. The thin-sheet equations are based on the lubrication theory, which assumes that the film thickness (normal to the surface) is much smaller than the characteristic length scale along the surface. Consequently, these equations break down at large contact angles, where the interface height violates the underlying thin-film assumption. We note that our LBM method solves the Navier-Stokes equations and is valid for the whole range of contact angles. In the case of a contact angle of $\theta=90^{\circ}$, we simulate the coalescence of two suspended spherical droplets in another liquid, and interestingly, we find that the theoretical prediction based on the thin-sheet equations underestimates the bridge growth - contrary to that at a contact angle of $\theta=59^{\circ}$.
@@ III.3 Two liquid lenses - 3D
In Fig. 8(a), we show the time evolution of the bridge height $h_{0}$ and radius $r_{0}$ for different contact angles $\theta=60^{\circ}$, $43^{\circ}$, respectively. The vertical height of the bridge grows faster with a larger contact angle, similar to what was observed in the $2D$ simulations (see Fig. 6(a)). Interestingly, the horizontal radius of the bridge grows at the same speed for both contact angles in the initial stage, which indicates that the bridge radius growth is not affected by the initial contact angle of the lenses, and the growth of the bridge radius is not a linear function of the growth of bridge height, in contrast to that observed in coalescing sessile droplets on a substrate. A possible explanation is that due to the no-slip boundary condition, the coalescence speed of sessile droplets on a substrate is relatively slow, giving sufficient time for the cross-section of the bridge to relax immediately to a spherical cap shape with a corresponding equilibrium contact angle. However, the coalescence speed of liquid lenses at a fluid-fluid interface is significantly faster, and the bridge height and radius grow in a non-coupled manner, resulting in a non-linear dependence of bridge radius and height growth. Although our simulations were limited to contact angles of $\theta=43^{\circ}$ and $60^{\circ}$ due to computational resource constraints, the initial growth of the bridge radius is expected to be independent of the contact angle across a wider range. This expectation is based on previous work, which shows that the growth of the width of coalescing lenses agrees with studies for freely suspended, respectively spherical droplets Scheel2023. If one considers a horizontal cross-section through the liquid lenses, their coalescence resembles that of freely suspended droplets with an effective contact angle of $90^{\circ}$. In this view, the contact angle observed in the vertical cross-section plays a negligible role in the initial stage.
@@ IV Conclusion
We numerically investigate the inertial coalescence of liquid lenses over a wide range of contact angles using the pseudopotential lattice Boltzmann method. While prior work has focused largely on liquid lenses with small contact angles and only addresses the initial stage of coalescence, our work extends the analysis to the coalescence of liquid lenses with large contact angles and the coalescence dynamics at later stages.
Our simulations in two dimensions successfully capture the self-similar dynamics of both the bridge and velocity profiles, showing quantitative agreement with experimental observations and a theoretical framework based on thin-sheet equations. Furthermore, our simulation results demonstrate that the thin-sheet equations are applicable to describe the quantitative behavior of bridge growth for small contact angles approximately up to $\theta<40^{\circ}$. In the three-dimensional case, we find out that the bridge height grows faster with a larger contact angle, while the growth of the bridge radius is independent of the contact angle at the initial stage. The contact angle of the cross-section of the bridge increases rapidly in the initial stage. It reaches a plateau at the intermediate stage, indicating a transition from a non-linear to a linear dependency between the growth of bridge height and bridge width. A theoretical model capable of predicting the transition’s critical conditions—such as the critical bridge radius or time—and describing the full bridge growth dynamics is not yet available. Developing such a model extends beyond the scope of the current work and represents a valuable direction for future research.
The pseudopotential lattice Boltzmann method has been widely applied to study binary fluid component systems Kruger2017; liu_multiphase_2016; xie_effect_2025, and our results demonstrate its applicability to ternary fluid systems, potentially inspiring its use for investigating printing and coating multi-component solutions for functional material synthesis in catalytic and electronic applications vinodh_recent_2024; steinberger_challenges_2024. For instance, in wet-on-wet inkjet printing, the final deposition pattern is affected by the competition between coalescence and evaporation timescales Wijshoff2018. Our findings provide an estimate of the coalescence timescale, thereby offering practical guidelines for optimizing drying conditions.